# OminiControl: Minimal and Universal Control for Diffusion Transformer

### v1

@@ 3.2 Image condition integration
where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$. This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2). rect addition approach.
@@
Comparative analysis shows that our approach achieves better results than the direct additive method, as demonstrated by the attention visualization in Figure 3. Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the di The effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
@@ 3.3 Adaptive position embedding
The integration of condition images requires careful consideration of positional information to ensure effective interaction between condition and target images. Previous methods often assume rigid spatial alignment between condition and output images, limiting their applicability to non-aligned tasks. Moreover, the relative positioning between condition and target tokens can significantly impact the model’s learning efficiency and generalization ability. Leveraging our unified sequence design where condition tokens are concatenated with other tokens, we have the flexibility to explore different positional encoding strategies.
In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information. For a 512$\times$512 input image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token receives a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$. This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the original image. However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)). Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
@@ 4.2 Main result
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on Canny-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [38], and MUSIQ [10] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [8] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
@@ 4.3 Empirical studies
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like canny-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
@@ 5 Conclusion
OmniControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OmniControl’s scalability and effectiveness in diffusion models.

### v2

@@ 3.2 Image condition integration
where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$. This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2).
@@
The comparative result shows that our approach achieves higher generation quality and better alignment with the conditions compared to the direct adding method, as illustrated in Figure 3. Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the direct adding approach. Besides, the effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
@@ 3.3 Adaptive position embedding
Our unified sequence design allows for flexible integration of condition image tokens, but this requires incorporating positional information to ensure effective interaction with noisy image tokens. The relative positioning of these tokens is critical, as it directly affects the model’s learning efficiency and generalization capability.
In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information. For a 512$\times$512 target image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token is assigned a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$. This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the noisy image. However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)). Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
@@ 4.2 Main result
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [38], and MUSIQ [10] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [8] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
@@ 4.3 Empirical studies
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like depth-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
@@ 5 Conclusion
OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OminiControl’s scalability and effectiveness in diffusion models.
6 Acknowledgment
We would like to acknowledge that the computational work involved in this research work is partially supported by NUS IT’s Research Computing group using grant numbers NUSREC-HPC-00001.